OpenStudy (anonymous):

find the x and y intercept of the equation 3y=6x -12

OpenStudy (bradely):

Find y intercept x=0 3y=6(0) -12 3y=-12 y=-4 Find x intercept y=0 3y=6x -12 0=6x-12 6x=12 x=2 Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

thanks bradly...can u give me the slope of this equation???

OpenStudy (bradely):

3y=6x -12 Divide each side by 3 3y/3=6x/3 -12/3 y=2x-4 Slope 2

OpenStudy (anonymous):

tahnks a lot...now i know much better about getting the intercepts and slope...:D